### v1

@@ 4.2 Solving Coom[x]in ASP
We represent the configuration as a tree such that each node represents a variable, and its root reflects the object to be configured. Variables can be either parts or attributes. Moreover, not all variables are necessarily included in the solution and an excluded variable renders all variables in its (possible) subtree excluded. Cardinalities are represented as constraints over sets of nodes in the tree.
@@ 4.2.1 Fact format
In Coom, a partonomy is established by defining structures and linking them through features. These features can also reference enumerations or numeric variables, potentially with cardinalities (as illustrated in Listing 5). We simplify this approach by treating both part and attribute variables uniformly and refer to the resulting structure as the configuration tree. For the TravelBike this structure, which serves as the basis for our fact format, is visualized in Figure 2. Here, nodes belonging to parts are highlighted in yellow and those belonging to attributes in green. The third bag of the carrier and the second bag of the frame are highlighted in a lighter color, meaning that these variables are undefined and not included in the solution. This automatically renders their subnodes undefined as well. Note that for the sake of readability variable names are abbreviated. Cardinalities of features are treated by grouping variables belonging to the same feature and with the same parent variable in sets (represented by dashed circles in the diagram).
Our refinement process generates all possible variables within the model and assigns them unique (human-readable) identifiers by leveraging the tree-like structure of the configuration. For example, the frame feature of the product in Listing 5 results in the variable "root.frame[0]". In this context, "root" refers to the object being configured, in our case, the TravelBike. Indices are required as feature cardinalities are potentially larger than 1. For instance, the bag feature of the Frame structure yields two variables "root.frame[0].bag[0]" and "root.frame[0].bag[1]" which can be read as “the first and second bag of the first frame”, respectively. For capturing each such variable in an object-centered representation, we include three facts standing for the type, index and parent of the variable.
@@
As before, a variable is associated with this attribute through predicate type/2, eg.  type("root.totalVolume[0]","totalVolume"). Currently, the only numeric attribute type is integer.
@@
The predicate function/3 comprises the function’s identifier, its type (eg.  count), and a set identifier. As above, the members of the set are declared via set/2.
@@ 4.2.2 Encoding
Listing 13 contains rules for the count and sum aggregate function as well as for the + operator. For all three rules we are making use of built-in clingo functionality.
@@
4.3 Solving Coom[x] with fclingo
An alternative Coom[x] encoding is given for solver fclingo (Cabalar et al., 2025) which is a prototype system for solving conditional linear constraints with integer variables in ASP. This system is a continuation of the lc2casp system (Cabalar et al., 2016) enhanced with conditional linear constraints given via the translation in (Cabalar et al., 2020). fclingo uses CASP solver clingcon (Banbara et al., 2017) as a backend, therefore being able to deal with large integer ranges and numerical calculations. The main difference, however, is that variables in fclingo can stay undefined while in clingcon all variables need to have a value assigned. As Coom allows for optional attributes, this is a crucial feature. The encoding for fclingo is identical in many parts to the clingo encoding from the previous section, except for rules which treat integer attributes and numerical constraints.
We start by giving a brief overview of the fclingo features we are utilizing for this adaptation and then proceed to highlight the differences between the two encodings. Our main objective with the fclingo encoding is to improve performance for numerical instances by overcoming clingo’s grounding bottleneck. This is achieved in fclingo by allowing for a special type of variables which are not subject to grounding. These are called integer variables and as their name suggests, can take values from the domain of the integers. Further, to be able to properly replace the necessary rules from our earlier encoding, we require that these integer variables can be used to build linear constraints while comprising a notion of undefinedness and the possibility to define defaults. Lastly, we also need to be able to exclude variables from a calculation if they do not fulfill certain conditions, ie. aggregate functions should allow for conditionality. fclingo provides us with all these features as we see in the following. Rules in fclingo are written in the same way as in clingo but additionally there are a few special theory atoms with which we can define integer variables and build up constraints over them. Here, we explain only those theory atoms that we are using in our encoding and refer the reader to Cabalar et al. (2025) for further documentation and examples. We proceed by showing snippets of the fclingo encoding while highlighting the differences from the clingo encoding from Section 4.2.2.
First, Line 16 in Listing 14 represents the fclingo replacement of the rule in Lines 15 and 16 in Listing 11 which is responsible for generating possible integer attribute values.
@@ 4.3 Solving Coom[x] with fclingo
Listing 14: Generation of integer attribute values in fclingo
Here, we make use of the fclingo atom &in which assigns a value from a range of integers to a variable. Note that variable X is not subject to grounding as it refers to the name of an integer variable. The operator =: specifies that this is a directional assignment, meaning that X only gets a value assigned if the lower and upper bound of the interval are defined. In this case, this would not be necessary as L and U are constants, however, the current syntax of fclingo only allows to use &in with directional assignments.
Due to the fact that integer variables in fclingo are freed from grounding (and unlike in the clingo encoding), it is possible to reason with unbounded ranges (cf. Line 17). When no range is specified, the &sum{X} = X atom defines variable X without constraining its range, thus effectively assigning any possible, integer value to it.
@@
Listing 15: Boolean constraint checking in fclingo
In Line 41 we make use of fclingo’s &df atom to check whether a variable is defined. Binary formulas are checked for satisfaction in the same way as before with the difference that we are using the fclingo atom &sum. Recall that since integer variables are treated in a special way in fclingo, they may only be accessed in the context of theory atoms (or constraints formed by them as in the case of variable X2).
Lastly, in Listing 16 we show rules for numerical calculations in fclingo which correspond to the rules from Listing 13 for clingo.
@@
Listing 16: Numerical calculations in fclingo
They represent calculations of count and sum aggregates, as well as for the + operator. While in the earlier encoding a different clingo functionality was used for each rule, here we use the fclingo atom &sum for all calculations. For the count function, we utilize the &sum atom with weight $1$ and for the + operator, we explicitly specify the two variables we want to add.
As all these &sum atoms appear in the heads of rules, they assign values to variables. Note, however, that here we are not using the directional assignment operator =: but just normal equality. In general, fclingo offers two distinct semantics: One, which is evoked by using the operator =: and which only assigns a value to the variable on the right side, if all variables on the left side are defined. And a second one, evoked by using normal equality which has no directionality and forces all variables occuring in the constraint to take on a value. In that case, if both variables in the constraint &sum{X} = Y were undefined, the constraint would force them to take any value as long as they are equal (which could possibly result in an infinite amount of solutions).
@@
We can solve a Coom instance using fclingo by adding the --solver fclingo (or -s fclingo for short) option on the command-line. By default, the CoomSuite uses clingo as solver. We are not showing the command-line output here as it coincides with the one from clingo.
@@ 4.4 Adding user input
All the Coom language fragments described so far equip the user with the ability to build up a configuration model which can be seen as a blueprint for all possible configurations. Another important concept in configuration is that of user requirements which allows the user to specify knowledge that needs to be included in the current configuration (Soininen et al., 1998). For example, the user could require the solution to include an optional component (eg. the basket of a bike) or to set the value for an attribute (eg. defining a minimum number of storage space in the TravelBike).
@@
Note that as as upper bounds of cardinalities are compiled away during the refinement process, it is not necessary to check for exceeding maximal cardinalities. Instead, this is implicitly handled by the first kind of warning.
@@
Independently of any warnings, the user input is processed by the encoding in Listing 18.
@@
However, before adding any user input to the final configuration, the encoding first verifies for each user_include/1 and user_value/2 that the user input is consistent with the configuration model (Lines 1-3). We omit here the rules for integer domains as the check is similar as for discrete domains.
@@ 4.5 Reasoning with unbounded cardinalities
10 small = ( 10 12 )
11 medium = ( 15 16 )
12 large = ( 25 20 )
13}
14behavior {
@@
In all earlier Coom language fragments a lower and upper bound for feature cardinalities had to be explicitly specified (or the cardinality defaulted to $1$ if no bounds were given). In practice, however, when modeling a configuration problem, it is often not possible to determine the exact (or even approximate) number of objects needed in advance. In this section we extend Coom[p] and Coom[x] to support unbounded cardinalities and call these Coom[p*] and Coom[x*], respectively. Then, we present a simple modification of our existing workflow to solve problems of such kind.
Listing 20 shows the CargoBike, a slightly modified version of the TravelBike example from Listing 5 such that the number of possible bags is now unknown. As in the TravelBike, the CargoBike has two numeric features totalVolume and requestedVolume where the former is computed by summing up the volume of all bags and the latter specifies a minimum for the former (computed) value. However, different from before, the feature bags has cardinality 0..* which stands for zero or more bags and is unbounded above. This can not be solved natively with the encoding presented in Section 4.2.
As a simple workaround, we set a maximum bound during the refinement phase and incrementally increase this bound until reaching a solution. We call this approach Incremental Bounds and proceed in the following manner (cf. Listing 3):
@@
    If no solution is found, increment the maximum upper bound by $k$ (typically $k=1$) and repeat from step 3.
Figure 3: Workflow for Incremental Bounds algorithm
While this approach is very simple, its advantages are that it is general (works with any combination of open cardinalities) and does not require any changes to the ASP encoding. Further, it provides a built-in minimization of the number of parts.
To solve a Coom instance with incremental bounds, we can use the --incremental-bounds option as shown in Listing 21. Here, we have set the value of requestedVolume to $60$ via a user input file. As the maximum volume of a single bag is $20$, the iteration is not able to find a solution until the maximum bound is set to $3$.
\$ coomsuite cargo-bike.coom -u input.coom --incremental-bounds -o coom
@@ 5 Solving Coom interactively
While in the previous sections the workflow assumed an input from the user only at the beginning of the process, in practice, users typically want to interact with the system continuously, eg. to experiment with different settings, explore the solution space or just to debug the configuration model. This is especially true for complex configuration problems where the number of solutions is large and difficult to envision by just looking at an input model (whether in textual or graphical form). In that sense, a user interface (UI) for solving Coom interactively is a valuable addition to the CoomSuite workbench.
For that purpse, we next present a prototypical UI which is generated and driven by ASP, more precisely by the clinguin system (Beiser et al., 2025). This system uses a simple design with dedicated predicates to define a UI and the behavior of user-triggered events, thereby greatly facilitating the specification of continuous user interactions with an ASP system. For details about the syntax and functionality of clinguin we refer the reader to (Beiser et al., 2025).
The UI presented in this section uses the Coom[x] encoding presented in Section 4.2 plus an additional clinguin encoding defining its layout, style and functionality. Since integration with constraint systems such as fclingo is currently limited in clinguin, the UI only works with clingo.
@@
We start by showcasing a simple interaction of the UI using the TravelBike example by presenting snippets of the UI encoding together with screenshots of the possible user interactions. The full UI encoding can be found in (Baumeister et al., 2025a). Then, we proceed to show the explanation features of the UI and how they are implemented in clinguin.
@@
With these auxiliary predicates, the first rule in Listing 24 creates a dropdown menu on the UI for each attribute and the second rule adds the text for this dropdown menu for when a value exists. We can see this when comparing the first and second screenshot in Figures 4 and 5 where the text in the dropdown menus of the color and frontWheel enumerations changes accordingly.
@@
The style of the dropdown menus is defined in Listing 25 where the first two rules set the style for inferred values. The first rules add the light text with low opacity for values that are inferred but not (yet) selected, eg. as for the frontWheel. Furthermore, the second rule removes the border of the dropdown menus of such attributes to better distinguish them.
Next, the third and fourth rule add the class "btn-secondary", which is the purple button style, for mandatory discrete attributes (stemming from enumerations in Coom) for which either a value has been selected or there are multiple options left. Note that for numeric values (not shown here) this works similarly but with the class "btn-primary", which is the blue button style.
@@
The rules in Listing 26 are responsible for providing the options of the dropdown menus that are still valid. The first rule creates an auxiliary predicate with all possible values which is subsequently used in the last three rules to create one dropdown menu item for each such value. Upon clicking on one of these values, a call is made to the solver to add the assumption of the corresponding value assignment, thus forcing the entailment of the given value and restricting the possible solutions (cf. Line 32).
27v_option((X,V),valid) :- i_attr(X,_), _any(value(X,V)), %%
28 not _clinguin_assume(value(X,_),true). %%
@@
Our approach to explanations for invalid values follows (Beiser et al., 2025). In a nutshell, this means that we use a specialized clinguin backend which provides a minimal unsatisfiable set (MUS) of assumptions whenever the encoding is unsatisfiable. In this way, we allow the user to select invalid values, and thus to let the problem become unsatisfiable, while subsequently using the information about faulty assumptions to highlight previously selected values in red. For our approach, this has the limitation that it only works for invalid selections of attribute values but not for the inclusion of an invalid part.
Additionally, we leverage the definition of constraints in the refined ASP instance to provide the user with a natural language explanation of why a certain value is invalid. To do this, we indicate to the explanation backend that the atoms of predicate constraint/2 should be treated as assumptions, and thus be included as part of the reasons for unsatisiablity. Moreover, we extend the ASP fact format for the configuration to include the predicate configuration_explanation/2 to store natural language explanations for Boolean constraints. For instance, the constraint !root.color[0]=Red||root.frontWheel[0].size[0]=20 has a natural language explanation: “If the color is red, then the size of the front wheel should be 20.”
Thanks to the generality of Boolean operators, we were able to automatically generate these explanations for all Boolean constraints using an LLM by prompting it with two simple examples. However, they can also be provided by the user. In fact, Coom offers a dedicated keyword for this which we plan to integrate in the workflow of the CoomSuite in the future. While it is desirable to have these kinds of explanations for all constraints, we limit ourselves to Boolean constraints for now and leave explanations for table constraints as future work as their treatment is more complex.
@@
Finally, leaving behind the topic of explanations, the last screenshot shows the addition of a bag to the configuration (cf. Figure 7) by means of clicking on the small icon containing a $+$. Since it was optional to add this part, it includes a red button to remove it. For the sake of brevity, we do not go into further detail here.
@@
As next steps, the user might decide to browse the different solutions, pick one and continue modifying it, or download it in the format of a Coom solution. The interested reader can try this out by following the instructions in (Baumeister et al., 2025a). Currently, this requires running clinguin separately but for future versions of the CoomSuite a direct integration is planned.

### v2

@@ 4.2 Solving Coom[x] in ASP
We represent the configuration as a tree such that each node represents a variable, and its root reflects the object to be configured. Variables can be either parts or attributes. Solutions are represented as subtrees of this configuration tree where values are assigned to any attribute variable included in the subtree. Thus, not all variables are necessarily included in the solution and an excluded variable renders all variables in its (possible) subtree excluded. Cardinalities are represented as constraints over sets of nodes in the tree.
@@ 4.2.1 Fact format
In Coom, a partonomy is established by defining structures and linking them through features. These features can also reference enumerations or numeric variables, potentially with cardinalities (as illustrated in Listing 5). We simplify this approach by treating both part and attribute variables uniformly and refer to the resulting structure as the configuration tree which can be considered an explicit representation of the Coom configuration product model from before.
For the TravelBike, this structure, which serves as the basis for our fact format, is visualized in Figure 2. Here, nodes belonging to parts are highlighted in yellow and those belonging to attributes in green. Consider an example solution where the third bag of the carrier and the second bag of the frame are undefined, thus not included in the solution. We represent this in Figure 2 by highlighting the corresponding variables in a lighter color. Note that undefined variables automatically render their subnodes undefined as well. Note that for the sake of readability variable names are abbreviated and value assignments are omitted. Cardinalities of features are treated by grouping variables belonging to the same feature and with the same parent variable in sets (represented by dashed circles in the diagram).
Our refinement process generates all possible variables within the model and assigns them unique (human-readable) identifiers by leveraging the tree-like structure of the configuration. For example, the frame feature of the product in Listing 5 results in the variable "root.frame[0]". In this context, "root" refers to the object being configured, in our case, the TravelBike. Indices are required as feature cardinalities are potentially larger than 1. For instance, the bag feature of the Frame structure yields two variables "root.frame[0].bag[0]" and "root.frame[0].bag[1]" which can be read as “the first and second bag of the first frame”, respectively. For capturing each such variable in an object-centered representation, we include three facts representing the type, index and parent of the variable.
@@
As before, a variable is associated with this attribute through predicate type/2, e.g., type("root.totalVolume[0]","totalVolume"). Currently, the only numeric attribute type is integer.
@@
The predicate function/3 comprises the function’s identifier, its type (e.g., count), and a set identifier. As above, the members of the set are declared via set/2.
@@ 4.2.2 Encoding
Listing 13 contains rules for the count and sum aggregate function as well as for the + operator. For all three rules we are making use of built-in clingo functionality. Note that whenever non-domain predicates (include/1 and value/2) occur inside an aggregate, assigning the result of that calculation to another variable possibly creates very large groundings. This is due to the fact that all possible value assignments have to be explicitly represented. However, in this encoding we opted for clarity and simplicity over performance and leave optimized versions for future work.
@@
4.3 Solving Coom[x] with flingo
An alternative Coom[x] encoding is given for solver flingo (Cabalar et al. 2025)555fclingo in earlier versions. which is a prototype system for solving conditional linear constraints with integer variables in ASP. This system is a continuation of the lc2casp system (Cabalar et al. 2016) enhanced with conditional linear constraints given via the translation in (Cabalar et al. 2020). flingo uses CASP solver clingcon (Banbara et al. 2017) as a backend, therefore being able to deal with large integer ranges and numerical calculations. The main difference, however, is that variables in flingo can stay undefined while in clingcon all variables need to have a value assigned. As Coom allows for optional attributes, this is a crucial feature. The encoding for flingo is identical in many parts to the clingo encoding from the previous section, except for rules which treat integer attributes and numerical constraints.
We start by giving a brief overview of the flingo features we are utilizing for this adaptation and then proceed to highlight the differences between the two encodings. Our main objective with the flingo encoding is to improve performance for numerical instances by overcoming clingo’s grounding bottleneck. This is achieved in flingo by allowing for a special type of variables which are not subject to grounding. These are called integer variables and as their name suggests, can take values from the domain of the integers. Importantly, they are being treated differently than propositional variables. Further, to be able to properly replace the necessary rules from our earlier encoding, we require that these integer variables can be used to build linear constraints while comprising a notion of undefinedness and the possibility to define defaults. Lastly, we also need to be able to exclude variables from a calculation if they do not fulfill certain conditions, i.e., aggregate functions should allow for conditionality. flingo provides us with all these features as we see in the following. Rules in flingo are written in the same way as in clingo but additionally there are a few special theory atoms through which we can define integer variables and build up constraints over them. Here, we explain only those theory atoms that we are using in our encoding and refer the reader to A and Cabalar et al. (2025) for further documentation and examples. We proceed by showing snippets of the flingo encoding while highlighting the differences from the clingo encoding from Section 4.2.2.
First, Line 16 in Listing 14 represents the flingo replacement of the rule in Lines 15 and 16 in Listing 11 which is responsible for generating possible integer attribute values.
@@ 4.3 Solving Coom[x] with flingo
Listing 14: Generation of integer attribute values in flingo
The head atom &in{L..U} =: X assigns a values from the range L..U to variable X. Note that since variable X appears inside the theory atom, it is declared as an integer variables, and therefore not subject to grounding. The operator =: specifies that this is an assignment, meaning that X only gets a value assigned if the lower and upper bound of the interval, L and U, are defined. In this case, this would not be necessary as both are constants, however, the current syntax of flingo only allows to use &in together with the assignment operator =:.
Due to the fact that integer variables in flingo are freed from grounding (and unlike in the clingo encoding), it is possible to reason with unbounded ranges (cf. Line 17). When no range is specified, the atom &sum{X} = X works similar to a choice rule in ASP in the sense that it defines variable X without constraining its range, thus effectively assigning any possible, integer value to it.
@@
Listing 15: Boolean constraint checking in flingo
In Line 41 we make use of flingo’s &df atom to check whether a variable is defined. Binary formulas are checked for satisfaction in the same way as before with the difference that we are using the flingo atom &sum. Recall that since integer variables are treated in a special way in flingo, they may only be accessed in the context of theory atoms (or constraints formed by them as in the case of variable X2).
Lastly, in Listing 16 we show rules for numerical calculations in flingo which correspond to the rules from Listing 13 for clingo.
@@
Listing 16: Numerical calculations in flingo
They represent calculations of count and sum aggregates, as well as for the + operator. While in the earlier encoding a different clingo functionality was used for each rule, here we use the flingo theory atom &sum for all calculations. For the count function, we utilize the &sum atom with weight $1$ and for the + operator, we explicitly specify the two variables we want to add. For example, the atom &sum{ 1,X : set(P,X), include(X) } = F assigns to variable F the number of variables X that are contained in set P and included in the solution.
As all these &sum atoms appear in the heads of rules, they assign values to variables. Note, however, that here we are not using the directional assignment operator =: but just normal equality. In general, flingo offers two distinct semantics: One, which is evoked by using the operator =:, only assigns a value to the variable on the right side, if all variables on the left side are defined. The second one, evoked by using normal equality, has no directionality and forces all variables occurring in the constraint to take on a value. In that case, if both variables in the constraint &sum{X} = Y were undefined, the constraint would force them to take any value as long as they are equal (which could possibly result in an infinite amount of solutions).
@@
We can solve a Coom instance using flingo by adding the --solver flingo (or -s flingo for short) option on the command-line. By default, the CoomSuite uses clingo as solver. We do not show the command-line output here as it coincides with the one from clingo.
@@ 4.4 Adding user input
All the Coom language fragments described so far equip the user with the ability to build up a configuration model which can be seen as a blueprint for all possible configurations. Another important concept in configuration is that of user requirements which allows the user to specify knowledge that needs to be included in the current configuration (Soininen et al. 1998). For example, the user could require the solution to include an optional component (e.g., the basket of a bike) or to set the value for an attribute (e.g., defining a minimum number of storage space in the TravelBike).
@@
Note that as upper bounds of cardinalities are compiled away during the refinement process, it is not necessary to check for exceeding maximal cardinalities. Instead, this is implicitly handled by the first kind of warning.
@@
Independently of any warnings, the user input is processed by the encoding in Listing 18. However, before adding any user input to the final configuration, the encoding first verifies for each user_include/1 and user_value/2 that the user input is consistent with the configuration model (Lines 1-3). We omit here the rules for integer domains as the check is similar as for discrete domains.
@@ 4.5 Reasoning with unbounded cardinalities
In all earlier Coom language fragments a lower and upper bound for feature cardinalities had to be explicitly specified (or the cardinality defaulted to $1..1$ if no bounds were given). In practice, however, when modeling a configuration problem, it is often not possible to determine the exact (or even approximate) number of objects needed in advance (Taupe et al. 2016). In this section we extend Coom[p] and Coom[x] to support unbounded cardinalities and call these Coom[p*] and Coom[x*], respectively. Then, we present a simple modification of our existing workflow to solve problems of such kind.
Listing 20 shows the CargoBike, a slightly modified version of the TravelBike example from Listing 5 such that the number of possible bags is now unknown.
@@
9 small = ( 12 )
10 large = ( 20 )
11}
12behavior {
@@
As in the TravelBike, the CargoBike has two numeric features totalVolume and requestedVolume where the former is computed by summing up the volume of all bags and the latter specifies a minimum for the former (computed) value. However, different from before, the feature bags has cardinality 0..*, which stands for zero or more bags and is unbounded above. This can not be solved natively with the encoding presented in Section 4.2.
Figure 3: Workflow for Incremental Bounds algorithm
As a simple workaround, we set a maximum bound during the refinement phase and check whether this leads to a solution. The CoomSuite currently offers two search strategies to determine this maximum bound: A linear and an exponential one (Francescutto et al. 2021). We call this approach Incremental Bounds and proceed in the following manner (cf. Figure 3):
@@
    If no solution is found, determine a new maximum bound according to the chosen strategy and repeat from step 3.
While the linear search strategy simply increases the maximum bound by a chosen value $k$ (typically $k=1$), the exponential one is inspired by Francescutto et al. (2021) and first doubles the current bound until finding a solution and then converges to the optimal bound.
While this approach is very simple, its advantages are that it is general (works with any combination of open cardinalities) and does not require any changes to the ASP encoding. Further, at least in an approximate way, it provides a built-in minimization on the number of parts. However, it does not (yet) make use of clingo’s multi-shot solving capabilities, and therefore grounding has to performed again on every iteration.
To solve a Coom instance with incremental bounds, we can use the --bounds option to specify the desired search strategy as shown in Listing 21. Here, we have chosen linear search and set the value of requestedVolume to $60$ via a user input file. As the maximum volume of a single bag is $20$, the iteration is not able to find a solution until the maximum bound is set to $3$.
\$ coomsuite cargo-bike.coom -u input.coom --bounds=linear -o coom
@@ 5 Solving Coom interactively
While in the previous sections the workflow assumed an input from the user only at the beginning of the process, in practice, users typically want to interact with the system continuously, e.g., to experiment with different settings, explore the solution space or just to debug the configuration model. This is especially true for complex configuration problems where the number of solutions is large and difficult to envision by just looking at an input model (whether in textual or graphical form). In that sense, a user interface (UI) for solving Coom interactively is a valuable addition to the CoomSuite workbench.
For that purpose, we next present a prototypical UI which is generated and driven by ASP, more precisely by the clinguin system (Beiser et al. 2025). This system uses a simple design with dedicated predicates to define a UI and the behavior of user-triggered events, thereby greatly facilitating the specification of continuous user interactions with an ASP system. For details about the syntax and functionality of clinguin we refer the reader to (Beiser et al. 2025).
The UI presented in this section uses the Coom[x] encoding presented in Section 4.2 plus an additional clinguin encoding defining its layout, style and functionality. Since integration with constraint systems such as flingo is currently limited in clinguin, the UI only works with clingo.
@@
We start by showcasing a simple interaction of the UI using the TravelBike example by presenting snippets of the UI encoding together with screenshots of the possible user interactions. The full UI encoding can be found in (Baumeister et al. 2025a). Then, we proceed to show the explanation features of the UI and how they are implemented in clinguin.
@@
With these auxiliary predicates, the first rule in Listing 24 creates a dropdown menu on the UI for each attribute and the second rule adds the text for this dropdown menu when a value exists. We can see this when comparing the first and second screenshot in Figures 4 and 5 where the text in the dropdown menus of the color and frontWheel enumerations changes accordingly.
@@
The style of the dropdown menus is defined in Listing 25 where the first two rules set the style for inferred values. The first rules add the light text with low opacity for values that are inferred but not (yet) selected, e.g., as for the frontWheel. Furthermore, the second rule removes the border of the dropdown menus of such attributes to better distinguish them.
Next, the third and fourth rule add the class "btn-secondary", which is the purple button style, for mandatory discrete attributes (stemming from enumerations in Coom) for which either a value has been selected or there are multiple options left. Note that for numeric values (not shown here) this works similarly but with the class "btn-primary", which is the blue button style. Selecting values in clinguin is done via adding assumptions to clingo for the corresponding atom and the predicate _clinguin_assume/2 is added by clinguin for every atom that has been assumed to be true or false.
@@
The rules in Listing 26 are responsible for providing the options of the dropdown menus that are still valid. The first rule creates an auxiliary predicate with all possible values which is subsequently used in the last three rules to create one dropdown menu item for each such value.
27v_option((X,V),valid) :- i_attr(X,_), _any(value(X,V)),
28 not _clinguin_assume(value(X,_),true).
@@
Upon clicking on one of these values, a call is made to the solver to add the assumption of the corresponding value assignment, thus forcing the entailment of the given value and restricting the possible solutions (cf. Line 32). The dedicated clinguin predicate _any/1 is the counterpart for _all/1 and specifies that an atom is contained in at least one stable model.
@@
Our approach to explanations for invalid values follows (Beiser et al. 2025). In a nutshell, this means that we use a specialized clinguin backend which provides a minimal unsatisfiable set (MUS) of assumptions whenever the encoding is unsatisfiable. In this way, we allow the user to select invalid values, and thus to let the problem become unsatisfiable, while subsequently using the information about faulty assumptions to highlight previously selected values in red. For our approach, this has the limitation that it only works for invalid selections of attribute values but not for the inclusion of an invalid part.
Additionally, we leverage the definition of constraints in the refined ASP instance to provide the user with a natural language explanation of why a certain value is invalid. To do this, we indicate to the explanation backend that the atoms of predicate constraint/2 should be treated as assumptions, and thus be included as part of the reasons for unsatifiablity. Moreover, we extend the ASP fact format for the configuration to include the predicate explanation/2 to store natural language explanations for Boolean constraints. For instance, the constraint !root.color[0]=Red||root.frontWheel[0].size[0]=20 has a natural language explanation: “If the color is red, then the size of the front wheel should be 20.”
Thanks to the generality of Boolean operators, we were able to automatically generate these explanations for all Boolean constraints using a Large Language Model (LLM) by prompting it with two simple examples. However, they can also be provided by the user. In fact, Coom offers a dedicated keyword for this which we plan to integrate in the workflow of the CoomSuite in the future. While it is desirable to have these kinds of explanations for all constraints, we limit ourselves to Boolean constraints for now and leave explanations for table constraints as future work as their treatment is more complex.
@@
Finally, leaving behind the topic of explanations, the last screenshot shows the addition of a bag to the configuration (cf. Figure 7) by means of clicking on the small icon containing a ++. Since it was optional to add this part, it includes a red button to remove it. For the sake of brevity, we do not go into further detail here.
@@
As next steps, the user might decide to browse the different solutions, pick one and continue modifying it, or download it in the format of a Coom solution. The interested reader can try this out by following the instructions in (Baumeister et al. 2025a). Currently, this requires running clinguin separately but for future versions of the CoomSuite a direct integration is planned.